OpenStudy (anonymous):

Use your graphing calculator to find all degree solutions in the interval 0° ≤ x < 360° for the following equation. (Enter your answers as a comma-separated list.) cos 3x = sqrt2/2 battery on my graphing calculator died, can anyone help?

OpenStudy (campbell_st):

well if you have some graphing software just graph y = cos(3x) and y = sqrt2/2 and find the points where they intersect this is where the 2 curves are equal.

OpenStudy (anonymous):

Don't have any access to anything that can graph atm..

OpenStudy (campbell_st):

well you ahve the net, so try wolfram alpha or fooplot

OpenStudy (anonymous):

Is it possible to do this without a graphing calculator?

OpenStudy (campbell_st):

sure... if you use \[\cos(3x) = \frac{\sqrt{2}}{2}\] this is an exact value... 45 degrees so \[3x = \cos^{-1} \frac{\sqrt{2}}{2}\] or 3x = 45 so x = 15 degrees well to make it easy here is a graph showing points of intersection hope it helps

OpenStudy (campbell_st):

oops sorry, just realised the graph is in radians.... so start at 15 and then just keep adding 90 degrees... to the previous...

OpenStudy (anonymous):

I've gotten as far as 3x=45, x=15 degrees but I don't think simply adding 90 degrees gets me the answers I need. If cos 3x=sqrt3/2=10, 110, 130,230,250,350 degrees which doesn't have any patterns such as adding 90 degrees to previous